### v1

# Minimisation of Laplacian eigenvalue with indefinite weight under inhomogeneous Robin boundary condition

Abstract
This paper explores a certain Laplacian eigenvalue optimisation problem with indefinite weight under inhomogeneous Robin boundary condition. The minimum principal eigenvalue is fully determined in one dimension by formulating the problem as a shape optimisation one. The result is verified numerically using a shooting method.
pacs:
MSC Classification]35P15, 35J25
1 Introduction
In this paper, we consider the following eigenvalue problem
$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&\ \text{in}&\ \Omega\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&\ \text{on}&\ \partial\Omega\end{aligned}\right$. (1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ with Lipschitz boundary, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, and the weight function $m$ is a bounded measurable function that changes sign in $\Omega$, i.e. $\Omega_{m}^{+}:=\{x\in\Omega;\ m(x)>0\}$ satisfies $|\Omega_{m}^{+}|\in(0,|\Omega|)$, and that
$1\leq m(x)\leq\kappa\ \text{a.e.}\ x\in\Omega$
for a constant $\kappa>0$. The inhomogeneous Robin boundary condition applies, with $\beta:\partial\Omega\to\mathbb{R}$ a piecewise continuous function.
According to [1], [2] and [3], $\lambda$ is called a principal eigenvalue of (1), if the associated eigenfunction $\phi$ is positive, where $\phi\in H^{1}(\Omega)$.
The investigation of Laplacian eigenvalues with an indefinite weight under Robin boundary conditions originates from studies of diffusive logistics models in a bounded domain $\Omega$ with suitable boundary. In [4], it is shown that the density $u(x,t)$ of the species at location $x$ and time $t$ satisfies the following equation:
$\begin{cases}u_{t}=\Delta u+\omega u[m(x)-u]\ \text{in}\ \Omega\times\mathbb{R}^{+},\\ \partial_{\mathbf{n}}u+\beta_{0}u=0\ \text{on}\ \partial\Omega\times\mathbb{R}^{+},\\ u(x,0)\geq 0,\ u(x,0)\not\equiv 0\ \text{in}\ \bar{\Omega},\end{cases}$ (2)
where $\omega>0$ is a certain parameter. The Robin boundary parameter $\beta_{0}$ represents the inhospitableness of $\partial\Omega$. If $\beta_{0}=0$, the boundary is a barrier; if $\beta_{0}=+\infty$, the boundary is hostile. $m$ represents the intrinsic growth rate of the species, changing sign in $\Omega$, so that $\int_{\Omega}m$ represents the total resources in the spatially heterogeneous environment. In [5], the relation between the minimal positive principal eigenvalue $\lambda=\lambda(m)$ of (1) and the parameter $\omega$ in (2) is discovered: if $\omega\leq\lambda(m)$, then $u(x,t)\to 0$ uniformly in $\bar{\Omega}$ as $t\to\infty$ for proper initial data, i.e. the species extinguish; if $\omega>\lambda(m)$, then $u(x,t)\to u^{*}(x)$ uniformly in $\bar{\Omega}$ as $t\to\infty$, where $u^{*}$ is the unique positive steady solution of (2) in $W^{2,q}(\Omega)$ for every $q>1$, i.e. the species survives. Consequently, it is natural to consider the problem of minimising such an eigenvalue.
In this paper we consider, more generally, the inhomogeneous Robin boundary condition, and we focus on the case $\beta\geq 0$ on $\partial\Omega$.
We first prove the existence and uniqueness of a positive principal eigenvalue following the method by [6].
Lemma 1 ([7], Theorem 11.4 and 11.10).
For $\phi\in H^{1}(\Omega)$, let
$Q_{\phi}(\lambda):={\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta\phi^{2}-\lambda\int_{\Omega}m\phi^{2}},$
and
$\mu(\lambda):=\inf_{\phi\in H^{1}(\Omega)}\frac{Q_{\phi}(\lambda)}{\int_{\Omega}\phi^{2}}.$
Then $\lambda$ is a principal eigenvalue of (1), if and only if $\mu(\lambda)=0$.
Lemma 2 ([6], Lemma 2).
If $\phi_{0}\in H^{1}(\Omega)$ is a minimiser of $Q_{\phi}(\lambda)$, i.e. $\mu(\lambda)=\frac{Q_{\phi_{0}}(\lambda)}{\int_{\Omega}\phi_{0}^{2}}$, then
$\mu^{\prime}(\lambda)=-\frac{\int_{\Omega}m\phi_{0}^{2}}{\int_{\Omega}\phi_{0}^{2}}.$
For any $\phi\in H^{1}(\Omega)$, the map $\lambda\mapsto Q_{\phi}(\lambda)$ is affine, hence concave. Since
$\mu(\lambda)=\inf\left\{Q_{\phi}(\lambda);\ \phi\in H^{1}(\Omega),\int_{\Omega}\phi^{2}=1\right\},$
the map $\mu:\lambda\mapsto\mu(\lambda)$ is also concave as the infimum of a collection of concave functions. In other words, $\mu^{\prime\prime}(\lambda)\leq 0$. Meanwhile, it is obvious that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$. Thus, $\mu=\mu(\lambda)$ increases until it attains its maximum, and decreases thereafter.
When $\beta>0$ on $\partial\Omega_{\beta}^{+}\subseteq\partial\Omega$ with $|\Omega_{\beta}^{+}|>0$, using the variational characterisation of the first Robin eigenvalue $\lambda_{1}(\beta)>0$, we have
$Q_{\phi}(0)\geq\lambda_{1}(\beta)\int_{\Omega}\phi^{2}>0$
whenever $\int_{\Omega}\phi^{2}>0$, hence $\mu(0)>0$, so $\mu$ has exactly two zeros, one positive and the other negative. In other words, (1) has a unique positive principal eigenvalue and a unique negative principal eigenvalue.
When $\beta\equiv 0$ a.e. on $\partial\Omega$, the situation is similar to Neumann boundary condition, and by taking the constant $\phi_{c}\equiv c$, we see
$Q_{\phi_{c}}(0)=0+c^{2}\int_{\partial\Omega}\beta=0,$
and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 2, we get:
  • if $\int_{\Omega}m<0$, then $\mu^{\prime}(0)>0$, and $\mu$ has a unique positive zero corresponding to the unique positive principal eigenvalue of (1);
  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)<0$, and $\mu$ has a unique negative zero corresponding to the unique negative principal eigenvalue of (1);
  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (1) has a unique principal eigenvalue 0.
It is also well-known (cf. [8], [9], [1] and [3]) that the eigenvalue can be expressed with a Rayleigh quotient. To sum up, we have:
Theorem 3.
If $\beta\geq 0$ on $\partial\Omega$ and $\beta>0$ on a subset of positive measure, the problem (1) admits a unique positive principal eigenvalue $\lambda$, depending on $m$. Moreover,
$\lambda(m)=\inf_{\phi\in\mathcal{S}(m)}\frac{\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta\phi^{2}}{\int_{\Omega}m\phi^{2}},$
where $\mathcal{S}(m):=\left\{\phi\in H^{1}(\Omega);\int_{\Omega}m\phi^{2}>0\right\}$.
Principal eigenvalues also admit monotonicity [^1] We remark that an attempt to derive strict monotonicity in [10], Theorem 2.4, fails, for a positive eigenfunction could vanish on the boundary. with respect to Robin parameter $\beta$.
Theorem 4.
Suppose $\lambda_{1},\lambda_{2}$ are principal eigenvalues of problem (1) with Robin parameters $\beta_{1}$, $\beta_{2}$, respectively. And denote by $\phi_{1}$ and $\phi_{2}$ the positive eigenfunctions associated with $\lambda_{1}$ and $\lambda_{2}$, respectively. Assuming $\int_{\Omega}m\phi_{1}^{2}>0$, if $\beta_{1}\geq\beta_{2}$, then $\lambda_{1}\geq\lambda_{2}$. If, in addition, we suppose that $\phi_{1}$ does not vanish on $\partial\Omega$ and $\beta_{1}>\beta_{2}$, then $\lambda_{1}>\lambda_{2}$.
Proof.
By Lemma 1,
$\int_{\Omega}|\nabla\phi_{1}|^{2}+\int_{\partial\Omega}\beta_{1}\phi_{1}^{2}-\lambda_{1}\int_{\Omega}m\phi_{1}^{2}$
$=0$
$=\inf_{\phi\in H^{1}(\Omega)}\left(\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta_{2}\phi^{2}-\lambda_{2}\int_{\Omega}m\phi^{2}\right)$
$\leq\int_{\Omega}|\nabla\phi_{1}|^{2}+\int_{\partial\Omega}\beta_{2}\phi_{1}^{2}-\lambda_{2}\int_{\Omega}m\phi_{1}^{2},$
so
$\lambda_{1}-\lambda_{2}\geq\frac{\int_{\partial\Omega}(\beta_{1}-\beta_{2})\phi_{1}^{2}}{\int_{\Omega}m\phi_{1}^{2}}\geq 0.$
We are interested in the minimisation of the positive principal eigenvalue $\lambda(m)$:
$\inf_{m\in\mathcal{M}_{m_{0},\kappa}}$ $\lambda(m),$ (3)
$\mathcal{M}_{m_{0},\kappa}$ $=\left\{m\in L^{\infty}(\Omega);\ -1\leq m\leq\kappa,\left|\Omega_{m}^{+}\right|>0,\int_{\Omega}m\leq-m_{0}|\Omega|\right\}$
Here $m_{0}$ is a real constant such that $m_{0}\in(-\kappa,1)$ if $\beta>0$ and $m_{0}\in(0,1)$ if $\beta=0$.
The following result is proven in [1] for Neumann boundary condition ($\beta\equiv 0$), but the extension to general Robin case is straightforward:
Theorem 5.
There exists some weight function $m^{*}\in\mathcal{M}_{m_{0},\kappa}$ such that the infimum of (3) is attained at $\lambda(m^{*})$. Moreover, there exists a measurable subset $E^{*}$ of $\Omega$ such that
$m^{*}=\kappa\chi_{E^{*}}-\chi_{\Omega\setminus E^{*}}\text{ a.e. }x\in\Omega,$ (4)
where $\chi_{U}$ denotes the characteristic function of a set $U$. In addition, the volume constraint is active at $m^{*}$:
$\int_{\Omega}m^{*}=-m_{0}|\Omega|.$
In other words, the optimal weight function $m^{*}$ satisfies a bang-bang distribution in $\Omega$ with respect to some optimal set $E^{*}$. The problem can therefore be formulated as a shape optimisation one:
$\inf_{E\in\mathcal{E}_{c,\kappa}}\lambda(E)$ $=\lambda(\kappa\chi_{E}-\chi_{\Omega\setminus E}),$ (5)
$\mathcal{E}_{c,\kappa}$ $=\{E\subseteq\Omega;E\ \text{is measurable},|E|\in(0,c|\Omega|]\}$
where $c:=\frac{1-m_{0}}{1+\kappa}$ and $c\in(0,1)$ if $\beta>0$ while $c\in\left(0,\frac{1}{1+\kappa}\right)$ if $\beta=0$. The optimal $E$ is expressed as a function of $\beta$, $\kappa$ and $c$.
In [2] and [3], this minimisation problem, formulated as a shape optimisation one, is thoroughly investigated under the usual homogeneous Robin boundary condition, with the 1-dimensional case fully solved. Unfortunately, for higher dimensional cases, few results are known. In [2], they studied this problem in a cylindrical domain with Robin boundary condition on the top and bottom and Neumann boundary condition on the lateral boundary. In [3], it is shown that when $\partial\Omega$ is connected, the optimal set is not always a ball. In [11], the 1-dimensional result is generalised onto spherical shells of any dimension.
2 The 1-dimensional case
In this section, we derive the complete solution of the problem (1) in one dimension.
Consider the problem (1) in $\Omega=(0,1)\subseteq\mathbb{R}$. The following result is shown in [3] for homogeneous Robin conditions, but its generalisation is straightforward.
Proposition 6.
If $\Omega=(0,1)$, then any optimal set $E^{*}$ for the problem (5) is an interval.
Now let the optimal set be the interval $E=(a,b)$ where $0<a<b<1$ and c=bac=b-a. We intend to find the optimal value of $a$ following the method by [2]. Let $\beta_{0}=\beta(0)\geq 0$ and $\beta_{1}=\beta(1)>0$. Rewrite the problem using $m=\kappa\chi_{E}-\chi_{\Omega\setminus E}$, and we have
$\left\{\begin{aligned} \phi^{\prime\prime}-\lambda\phi&=0&\ \text{in}&\ (0,a),\\ \phi^{\prime\prime}+\lambda\kappa\phi&=0&\ \text{in}&\ (a,b),\\ \phi^{\prime\prime}-\lambda\phi&=0&\ \text{in}&\ (b,1),\\ \phi^{\prime}(0)-\beta_{0}\phi(0)&=0,\\ \phi^{\prime}(1)+\beta_{1}\phi(1)&=0,\\ \phi(a)^{+}-\phi(a)^{-}&=0,\\ \phi(b)^{+}-\phi(b)^{-}&=0,\\ \phi^{\prime}(a)^{+}-\phi^{\prime}(a)^{-}&=0,\\ \phi^{\prime}(b)^{+}-\phi^{\prime}(b)^{-}&=0.\end{aligned}\right$. (6)
Since $\lambda>0$, $\phi>0$, we may assume
$\phi(x)=\left\{\begin{aligned} C_{1}\cosh\left(\sqrt{\lambda}(x-a)\right)+C_{2}\sinh\left(\sqrt{\lambda}(x-a)\right)\ \text{in}\ (0,a),\\ C_{3}\cos\left(\sqrt{\lambda\kappa}(x-a)\right)+C_{4}\sin\left(\sqrt{\lambda\kappa}(x-a)\right)\ \text{in}\ (a,b),\\ C_{5}\cosh\left(\sqrt{\lambda}(x-b)\right)+C_{6}\sinh\left(\sqrt{\lambda}(x-b)\right)\ \text{in}\ (b,1),\end{aligned}\right$. (7)
where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. Now we simplify the boundary and transmission conditions in (6). By organising the coefficients of $C_{i}$ ($1\leq i\leq 6$), we get
$\left\{\begin{aligned} &\left(-\sqrt{\lambda}\sinh(a\sqrt{\lambda})-\beta_{0}\cosh(a\sqrt{\lambda})\right)C_{1}+\left(\sqrt{\lambda}\cosh(a\sqrt{\lambda})+\beta_{0}\sinh(a\sqrt{\lambda})\right)C_{2}=0\\ &\left(\sqrt{\lambda}\sinh(\sqrt{\lambda}(1-b))+\beta_{1}\cosh(\sqrt{\lambda}(1-b))\right)C_{5}\\ &\ \ \ +\left(\sqrt{\lambda}\cosh(\sqrt{\lambda}(1-b))+\beta_{1}\sinh(\sqrt{\lambda}(1-b))\right)C_{6}=0\\ &C_{3}-C_{1}=0\\ &C_{5}-C_{3}\cos(\sqrt{\lambda\kappa}c)-C_{4}\sin(\sqrt{\lambda\kappa}c)=0\\ &\sqrt{\kappa}C_{4}-C_{2}=0\\ &C_{6}+C_{3}\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)-C_{4}\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)=0\end{aligned}\right$. (8)
Using the last 4 equations, we can express $C_{1},C_{2},C_{5},C_{6}$ in terms of $C_{3}$ and $C_{4}$, and then we obtain two linear equations of $C_{3}$ and $C_{4}$ only:
$\left\{\begin{aligned} &\left(-\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(a)-\beta_{0}{\mathfrak{c}}_{\lambda}(a)\right)C_{3}+\sqrt{\kappa}\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(a)+\beta_{0}{\mathfrak{s}}_{\lambda}(a)\right)C_{4}=0\\ &\left[\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)-\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]C_{3}\\ &+\left[\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)+\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]C_{4}=0\end{aligned}\right$. (9)
where ${\mathfrak{s}}_{\lambda}(a):=\sinh(\sqrt{\lambda}a)$, ${\mathfrak{c}}_{\lambda}(a):=\cosh(\sqrt{\lambda}a)$, ${\mathfrak{s}}_{\lambda}(b):=\sinh(\sqrt{\lambda}(1-b))$ and ${\mathfrak{c}}_{\lambda}(b):=\cosh(\sqrt{\lambda}(1-b))$. Considered as a $2\times 2$ matrix applied to the column vector $(C_{3},C_{4})^{t}$, they have a non-trivial solution if and only if the determinant of the matrix vanishes, i.e.
$\left(-\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(a)-\beta_{0}{\mathfrak{c}}_{\lambda}(a)\right)\left[\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ +\left.\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]\\ =\sqrt{\kappa}\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(a)+\beta_{0}\sinh(a\sqrt{\lambda})\right)\left[\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ -\left.\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right].$ (10)
First we order the term and get
$\cosh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-(\beta_{0}+\beta_{1})\sqrt{\lambda\kappa}\cos(\sqrt{\lambda\kappa}c)+(\lambda\kappa-\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{0}-\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right]\\ +\sinh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{1}-\beta_{0})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\beta_{0}\beta_{1}\kappa-\lambda)\sin(\sqrt{\lambda\kappa}c)\right)\right]=0.$ (11)
Let us first verify (11) in two extreme cases. When $\beta_{0}=\beta_{1}=0$, the problem is reduced to Neumann boundary condition, and we have obtained the characteristic equation [^2] We remark that the characteristic equations (18) and (20) obtained in [2] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
$\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-\tanh(\lambda^{\frac{1}{2}}a)}{1+\kappa^{-\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)},$ (12)
which reduces, when $a=0$, to the equation (3.3) in [1]:
$\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (13)
On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem is reduced to Dirichlet boundary condition, we similarly get the characteristic equation
$\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{-\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)+\tanh(\lambda^{\frac{1}{2}}a)}{\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-1},$ (14)
which reduces, when $a=0$, to the equation (21) in [2]:
$\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=-\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (15)
Now we use the product-to-sum formulae for hyperbolic functions to simplify it. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
$0=(\kappa+1)(\lambda-\beta_{0}\beta_{1})\cosh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +(\kappa+1)(\beta_{0}-\beta_{1})\sqrt{\lambda}\sinh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +\cosh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\lambda+\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)\right)\\ +\sinh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\beta_{0}+\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\kappa}(\beta_{0}\beta_{1}+\lambda)\cos(\sqrt{\lambda\kappa}c)\right).$ (16)
We express the right-hand side of (16) as a function $f(a,\beta_{0},\beta_{1},\kappa,\lambda)$, using $b=a+c$:
$f(a,\beta_{0},\beta_{1},\kappa,\lambda)=(\kappa+1)(\lambda-\beta_{0}\beta_{1})\cosh(\sqrt{\lambda}(2a+c-1))\sin(\sqrt{\lambda\kappa}c)\\ +(\kappa+1)(\beta_{0}-\beta_{1})\sqrt{\lambda}\sinh(\sqrt{\lambda}(2a+c-1))\sin(\sqrt{\lambda\kappa}c)\\ +\cosh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\lambda+\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)\right)\\ +\sinh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\beta_{0}+\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\kappa}(\beta_{0}\beta_{1}+\lambda)\cos(\sqrt{\lambda\kappa}c)\right).$ (17)
Taking partial derivative with respect to $a$ gives
$\partial_{a}f(a,\beta_{0},\beta_{1},\kappa,\lambda)=2(\kappa+1)\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\cosh(\sqrt{\lambda}(2a+c-1))\\ \cdot\left[(\lambda-\beta_{0}\beta_{1})\tanh\left(2\sqrt{\lambda}\left(a-\frac{1-c}{2}\right)\right)+\sqrt{\lambda}(\beta_{0}-\beta_{1})\right].$ (18)
Since we are concerned with the first eigenvalue only, we assume $0<\lambda<\frac{\pi^{2}}{4c^{2}\kappa}$, so $\sin(\sqrt{\lambda\kappa}c)>0$ and $\cos(\sqrt{\lambda\kappa}c)>0$. Therefore, the sign of $\partial_{a}f$ depends on that of
$g(a,\beta_{0},\beta_{1},\lambda):=(\lambda-\beta_{0}\beta_{1})\tanh\left(2\sqrt{\lambda}\left(a-\frac{1-c}{2}\right)\right)+\sqrt{\lambda}(\beta_{0}-\beta_{1}).$ (19)
Lemma 7.
For any $a\in[0,1-c]$, $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\not\equiv 0$, where $\kappa>0$.
Proof.
Write $\partial_{\beta_{1}}f=A(a)\,\beta_{0}+B(a)$, and assume that for some $a\in[0,1-c]$ we have $A(a)=B(a)=0$, where $A,B$ are defined in (37). Then $A(a)=B(a)=0$ gives
$(\kappa+1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(2a+c-1)\big)$ = $=-\,2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)$
$\quad+(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big),$ (20)
$(\kappa+1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(2a+c-1)\big)$ $=(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)$
$\quad-2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big).$ (21)
Now square both (20) and (21) and subtract them. Using the identity $\cosh^{2}\!\big(\sqrt{\lambda}(2a+c-1)\big)-\sinh^{2}\!\big(\sqrt{\lambda}(2a+c-1)\big)=1$, we obtain
$(\kappa+1)^{2}\sin^{2}\!\big(c\sqrt{\kappa\lambda}\big)=P^{2}-Q^{2},$ (22)
where
$P$ = $=-\,2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big),$ (23)
$Q$ $=(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)-2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big).$ (24)
Expanding $P^{2}-Q^{2}$,
$P^{2}-Q^{2}=\left[4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\sinh^{2}(\sqrt{\lambda}(1-c))+(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})\cosh^{2}(\sqrt{\lambda}(1-c))\right.\\ \left.-4\sqrt{\kappa}(\kappa-1)\cos(c\sqrt{\kappa\lambda})\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))\sinh(\sqrt{\lambda}(1-c))\right]\\ -\left[(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})\sinh^{2}(\sqrt{\lambda}(1-c))+4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\cosh^{2}(\sqrt{\lambda}(1-c))\right.\\ -\left.-4\sqrt{\kappa}(\kappa-1)\cos(c\sqrt{\kappa\lambda})\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))\sinh(\sqrt{\lambda}(1-c))\right]\\ =(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})-4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\\ =(\kappa+1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})-4\kappa,$ (25)
and comparing (22) and (25) we get $\kappa=0$, a contradiction to the assumption that $\kappa>0$. ∎
Lemma 8.
Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where
$c^{*}=\frac{1}{1+\frac{2\sqrt{\kappa}}{\pi}\ln\left(\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}\right)}.$ (26)
Then the function
$h(c,\kappa,\lambda)=\frac{(\kappa-1)\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))-2\sqrt{\kappa}\cos(c\sqrt{\kappa\lambda})\sinh(\sqrt{\lambda}(1-c))}{(\kappa+1)\sin(c\sqrt{\kappa\lambda})}$ (27)
satisfies $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$.
Proof.
Since $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$, $\sin(c\sqrt{\kappa\lambda})>0$ and $\cos(c\sqrt{\kappa\lambda})>0$, so we have
$h=\frac{\kappa-1}{\kappa+1}\cosh(\sqrt{\lambda}(1-c))-\frac{2\sqrt{\kappa}}{\kappa+1}\frac{\sinh(\sqrt{\lambda}(1-c))}{\tan(c\sqrt{\kappa\lambda})}.$ (28)
Let $t=c\sqrt{\kappa\lambda}\in\left(0,\frac{\pi}{2}\right)$ and $\theta=\sqrt{\lambda}(1-c)\in\left(0,\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)$, then
$h=h(\kappa,t,\theta)=\frac{\kappa-1}{\kappa+1}\cosh(\theta)-\frac{2\sqrt{\kappa}}{\kappa+1}\frac{\sinh(\theta)}{\tan(t)}<\frac{\kappa-1}{\kappa+1}\cosh(\theta)<\frac{\kappa-1}{\kappa+1}\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right).$ (29)
Therefore, to ensure $h<1$, it suffices that
$\frac{\kappa-1}{\kappa+1}\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)<1.$ (30)
Clearly, if $\kappa\leq 1$, (30) is satisfied automatically; if $\kappa>1$, (30) holds if and only if
$c>\frac{1}{1+\frac{2\sqrt{\kappa}}{\pi}\ln\left(\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}\right)}.$ (31)
We now prove:
Proposition 9.
Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is given in (26). Then there exists some $\beta_{0}^{*}=\beta_{0}^{*}(a)$ (given in (38)), such that $g(a,\beta_{0},\beta_{1},\lambda)$ and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)$ have opposite signs if $\beta_{0}>\beta_{0}^{*}$, and they share the same sign if $\beta_{0}<\beta_{0}^{*}$.
Proof.
We consider $\lambda=\lambda(a,\beta_{0},\beta_{1},\kappa)$ as implicitly defined by $f(a,\beta_{0},\beta_{1},\kappa,\lambda)=0$. Differentiate by $a$, and we get
$\partial_{\lambda}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)=-\partial_{a}f(a,\beta_{0},\beta_{1},\kappa,\lambda).$ (32)
Similarly, differentiate by $\beta_{1}$ and we get
$\partial_{\lambda}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)=-\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda).$ (33)
It is shown in Theorem 4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
$\partial_{\beta_{1}}f>0\Rightarrow\partial_{\lambda}f<0\Rightarrow\partial_{a}f\cdot\partial_{a}\lambda>0\Leftrightarrow g\cdot\partial_{a}\lambda>0,$ (34)
$\partial_{\beta_{1}}f<0\Rightarrow\partial_{\lambda}f>0\Rightarrow\partial_{a}f\cdot\partial_{a}\lambda<0\Leftrightarrow g\cdot\partial_{a}\lambda<0,$ (35)
$\partial_{\beta_{1}}f=0\Rightarrow\partial_{\lambda}f=0\Rightarrow\partial_{a}f=0\Leftrightarrow g=0.$ (36)
Notice that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)=A(a)\beta_{0}+B(a)$ is a linear function with respect to $\beta_{0}$, where
$A(a)=-2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(2a+c-1)\big),\\[5.69054pt] B(a)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(2a+c-1)\big).$ (37)
We prove in Lemma 7 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
$\beta_{0}=\beta_{0}^{*}(a):=-\frac{B(a)}{A(a)}.$ (38)
Moreover, when $\beta_{0}>\beta_{0}^{*}(a)$ or $\beta_{0}<\beta_{0}^{*}(a)$, the sign of $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ does not change, i.e. negative on one side and positive on the other. We now decide the monotonicity of $\partial_{\beta_{1}}f$ depending on $a$. The condition $A(a)>0$ is equivalent to
$\cosh(\sqrt{\lambda}(2a+c-1))<\frac{(\kappa-1)\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))-2\sqrt{\kappa}\cos(c\sqrt{\kappa\lambda})\sinh(\sqrt{\lambda}(1-c))}{(\kappa+1)\sin(c\sqrt{\kappa\lambda})}.$ (39)
Denote by $h=h(c,\kappa,\lambda)$ the expression on the right-hand side of (39). We show in Lemma 8 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$. In other words, $\partial_{\beta_{1}}f<0$ if $\beta_{0}>\beta_{0}^{*}$ and $\partial_{\beta_{1}}f>0$ if $\beta_{0}<\beta_{0}^{*}$. ∎
Lemma 10.
Denote by
$\tilde{\beta_{0}^{*}}(c,\kappa):=\frac{\sqrt{\kappa}\pi}{c}\frac{\sinh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)}{\kappa+1-(\kappa-1)\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)},$ (40)
where $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, and $c^{*}$ is as in (26). Then $\beta_{0}^{*}(a)$ defined in (38) has an upper bound $\beta_{0}^{*}\leq\tilde{\beta_{0}^{*}}$, not depending on $a$ or $\lambda$.
Proof.
Notice that
$A(a)\geq 2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ +(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\\ \geq-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)>0,$ (41)
where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 8. Meanwhile,
$B(a)\leq B(0)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ \leq 2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)$ (42)
hence we have
$\beta_{0}^{*}$ $\leq\frac{2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)}{-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)}$ (43)
$=\frac{2\kappa\sqrt{\lambda}\sinh\big(\sqrt{\lambda}(1-c)\big)}{(\kappa+1)-(\kappa-1)\cosh\big(\sqrt{\lambda}(1-c)\big)}.$ (44)
Notice that this upper bound is monotone increasing when $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$, hence
$\beta_{0}^{*}\leq\frac{\sqrt{\kappa}\pi}{c}\frac{\sinh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)}{\kappa+1-(\kappa-1)\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)},$ (45)
where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 8. ∎
Now we are ready to present the complete result on the 1-dimensional interval.
Theorem 11.
Consider problem (5) in $\Omega=(0,1)\subseteq\mathbb{R}^{1}$ with $\beta_{0}>\tilde{\beta_{0}^{*}}$, and $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is as in (26) and $\tilde{\beta_{0}^{*}}$ is as in (40). Then we have the following results. When $\beta_{0}\beta_{1}=\lambda$,
  • if $\beta_{0}>\beta_{1}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)<0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c;
  • if $\beta_{0}<\beta_{1}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $\beta_{0}=\beta_{1}=\lambda^{\frac{1}{2}}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)=0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for any $a\in[0,1-c]$.
Otherwise, denote by
$a^{*}=\frac{1}{2\sqrt{\lambda}}\tanh^{-1}\left(\frac{\sqrt{\lambda}(\beta_{1}-\beta_{0})}{\lambda-\beta_{0}\beta_{1}}\right)+\frac{1-c}{2}.$ (46)
When $\beta_{0}\beta_{1}>\lambda$,
  • if $\beta_{0}-\beta_{1}<-\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $|\beta_{0}-\beta_{1}|\leq\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in\left(0,a^{*}\right)$, and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in\left(a^{*},1-c\right)$, so the minimum eigenvalue is attained at $a=a^{*}$;
  • if $\beta_{0}-\beta_{1}>\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c.
When $\beta_{0}\beta_{1}<\lambda$,
  • if $\beta_{0}-\beta_{1}<-\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $|\beta_{0}-\beta_{1}|\leq\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,a^{*})$, and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(a^{*},1-c)$, so the minimum eigenvalue is attained at either $a=0$ or a=1ca=1-c;
  • if $\beta_{0}-\beta_{1}>\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c.
Proof.
The minimum of $\lambda$ can only be assumed at $a=0$, a=1ca=1-c or a possible internal minimal point, where $\partial_{a}\lambda=0$. By (32), a necessary condition satisfied by such an internal minimal point is $\partial_{a}f=0$. By (18) and (19), this happens if and only if $g=0$. First, suppose $\lambda=\beta_{0}\beta_{1}$. Then we see that if $\beta_{0}>\beta_{1}$, then $g>0$ for $a\in(0,1-c)$; if $\beta_{0}<\beta_{1}$, then $g<0$ for $a\in(0,1-c)$; and if $\beta_{0}=\beta_{1}$, then $g\equiv 0$ for $a\in(0,1-c)$.
Now we consider the case $\beta_{0}\beta_{1}>\lambda$. From (19), it is obvious that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
$\beta_{0}-\beta_{1}>\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (47)
and $a^{*}<0$ if and only if
$\beta_{0}-\beta_{1}<-\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (48)
so we have the following cases: if (48) is satisfied, then $g<0$ for $a\in(0,1-c)$; if (47) is satisfied, then $g>0$ for $a\in(0,1-c)$; and if
$\beta_{0}-\beta_{1}\leq\left|\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c))\right|,$ (49)
then $g>0$ for $a\in(0,a^{*})$ and $g<0$ for $a\in(a^{*},1-c)$.
Similarly we consider the case $\beta_{0}\beta_{1}<\lambda$. From (19), it is obvious that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
$\beta_{0}-\beta_{1}<-\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (50)
and $a^{*}<0$ if and only if
$\beta_{0}-\beta_{1}>\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (51)
so we have the following cases: if (50) is satisfied, then $g<0$ for $a\in(0,1-c)$; if (51) is satisfied, then $g>0$ for $a\in(0,1-c)$; and if
$\beta_{0}-\beta_{1}\leq\left|\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c))\right|,$ (52)
then $g<0$ for $a\in(0,a^{*})$ and $g>0$ for $a\in(a^{*},1-c)$.
This completes the proof together with Proposition 9. ∎
Remark on homogeneous Robin boundary
In [2], when studying the same problem under homogeneous Robin boundary ($\beta_{0}=\beta_{1}=\beta$), it is possible to reduce the condition $g=0$ to $\lambda=\beta_{0}\beta_{1}$, which could be plugged into $f=0$ to eliminate the terms involving $a$, which yields an equation of $\beta$, whose smallest positive solution can be expressed explicitly as $\beta^{*}=\beta^{*}(c,\kappa)$. Hence the final result in one dimension is given in cases where the criteria do not involve $\lambda$ which is a priori unknown. However, under our general inhomogeneous Robin boundary conditions, this is unfortunately impossible, since the condition $g=0$ does not help much in eliminating $a$ from the equation $f=0$ (cf. (19)) and (17). Nevertheless, in the next section we shall see that the criteria in Theorem 11 involving $\lambda$ could still be tested and thus the theorem itself be verified numerically.
3 Transfer–matrix shooting
In this section, we use a classical shooting method to numerically verify Theorem 11 for $\beta_{0}\beta_{1}\neq\lambda$.
For given $(c,\kappa,\beta_{0},\beta_{1})$ and a placement $a\in[0,1-c]$ of the favourable subinterval $(a,a+c)$, the equation
$\begin{cases}u^{\prime\prime}+\lambda\,m(x)\,u=0,\\ u^{\prime}(0)-\beta_{0}u(0)=0,\\ u^{\prime}(1)+\beta_{1}u(1)=0,\end{cases}$ (53)
with
$m(x)=\begin{cases}\kappa,&\text{ if }x\in(a,a+c),\\ -1,&\text{ otherwise,}\end{cases}$ (54)
admits explicit fundamental solutions on each subinterval:
$\begin{cases}u(x)=\rho\cosh(\sqrt{\lambda}\,x)+\sigma\sinh(\sqrt{\lambda}\,x),&x\in(0,a)\cup(a+c,1),\\ u(x)=\mu\cos(\sqrt{\lambda\kappa}\,x)+\nu\sin(\sqrt{\lambda\kappa}\,x),&x\in(a,a+c).\end{cases}$
Writing the state vector as $w=(u,u^{\prime})^{\top}$, then
$w^{\prime}(x)=(u(x)^{\prime},u(x)^{\prime\prime})^{\top}=\begin{pmatrix}0&1\\ -\lambda m(x)&0\end{pmatrix}w(x)=:Qw(x),$ (55)
where $m(x)\in\{-1,\kappa\}$ is piecewise constant. Now we integrate within an interval $x_{0},x]$ where $m(x)$ is constant, and get
$w(x)=e^{(x-x_{0})Q}w(x_{0}).$ (56)
The continuity of $u$ and $u^{\prime}$ at $x=a$ and $x=a+c$ thus yields
$w(1)=e^{(1-a-c)Q_{-1}}e^{cQ_{\kappa}}e^{aQ_{-1}}w(0),$ (57)
where for any $m\in\mathbb{R}$, $Q_{m}:=\begin{pmatrix}0&1\\ -\lambda m&0\end{pmatrix}$. Now that (53) is linear homogeneous with respect to $u$, we assume, without loss of generality that $u(0)=1$, then the Robin boundary condition gives $w(0)=(1,\beta_{0})^{\top}$ and $e_{2}w(1)+\beta_{1}e_{1}w(1)=0$, where $e_{1}=(1,0)$ and $e_{2}=(0,1)$. We now apply (57) and get
$R(a,\lambda):=e_{2}M(a,\lambda)\,(1,\beta_{0})^{\top}+\beta_{1}\,e_{1}M(a,\lambda)\,(1,\beta_{0})^{\top}=0,$ (58)
where $M(a,\lambda):=e^{(1-a-c)Q_{-1}}e^{cQ_{\kappa}}e^{aQ_{-1}}$. The principal eigenvalue we are interested in is interpreted as the smallest positive root of (58). We search for such roots on the window $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ using bisection on subintervals bracketed by a uniform scan in $\lambda$.
First, we pick a safe bound for $\lambda$ by setting
$\lambda_{\min}:=\max\left\{10^{-12},10^{-6}\frac{\pi^{2}}{4c^{2}\kappa}\right\},\ \ \ \lambda_{\max}:=(1-10^{-9})\frac{\pi^{2}}{4c^{2}\kappa},$ (59)
and take a uniform grid of sample $\lambda$-values:
$\lambda_{j}=\lambda_{\min}+j\frac{\lambda_{\max}-\lambda_{\min}}{N_{\lambda}},\ 0\leq j\leq N_{\lambda},$ (60)
where $N_{\lambda}$ is the number of subintervals we choose. Then a root of (58) must lie in a subinterval $(\lambda_{j},\lambda_{j+1})$ satisfying
$R(a,\lambda_{j})R(a,\lambda_{j+1})<0.$ (61)
If no such subinterval exists, we must refine the grid by increasing $N_{\lambda}$. Alternatively, we could lower $\lambda_{\min}$ to make it closer to 0.
Once a scan finishes with at least one subinterval where $R(a,\lambda)$ changes sign, we take the leftmost such subinterval $(\lambda_{j},\lambda_{j+1})$ and approach the eigenvalue $\lambda\in(\lambda_{j},\lambda_{j+1})$ by bisecting this subinterval. In other words, we check the sign of $\frac{\lambda_{j}+\lambda_{j+1}}{2}$ and refine the subinterval by its subinterval where $R(a,\lambda)$ changes sign, repeatedly until the length of the subinterval is smaller than a chosen tolerance, and output the eigenvalue as the midpoint of this final subinterval.
Finally, such numerical computation is done for a finite grid of $a$-values in the interval [0,1c][0,1-c], with a chosen reasonable resolution $N_{a}$, in order to produce the curve $a\mapsto\lambda(a)$ and to determine where the eigenvalue is minimised.
To compare the numerical result with Theorem 11 (non-degenerate cases $\beta_{0}\beta_{1}\neq\lambda$ only), we run the computation for a range of values for $(\beta_{0},\beta_{1})$ and for each pair of the Robin parameters, we verify the criteria in Theorem 11 for all computed eigenvalues $\lambda(a_{j})$, where $\{a_{j}\}_{j}$ are the grid of $a$-values in the interval [0,1c][0,1-c]. If a certain criterion in Theorem 11 is satisfied by all such $\lambda(a_{j})$, we classify the pair $(\beta_{0},\beta_{1})$ in the case corresponding to the valid criterion in Theorem 11, otherwise we do not classify this pair. For all those pairs of Robin parameters that are classified in a certain case of Theorem 11, we compare the predicted location for the minimum eigenvalue and the numerical result, and thus verify the correctness of Theorem 11.
Choice of parameters
We fix $N_{a}=81$, $N_{\lambda}=900$, and the tolerance $\texttt{tol}=10^{-10}$ of error. For Robin parameters, we sweep a rectangular grid $(\beta_{0},\beta_{1})\in[\beta_{\min},\beta_{\max}]^{2}$, with default $\beta_{\min}=0.2$, $\beta_{\max}=8.0$, and we take $n_{\beta}=10$ so that $n_{\beta}^{2}=100$ pairs are tested. The upper bound $\beta_{\max}$ is chosen to ensure that $\beta_{0}\beta_{1}-\lambda$ changes sign for the chosen $\kappa=2.0$ and $c=0.3$. In general, if some case from Theorem 11 is not found in a given $\beta$–grid, it is appropriate to widen $\beta_{\min},\beta_{\max}]$ or increase $n_{\beta}$.
Outputs
Our script produces:
  1. 1.
    a CSV file collecting all tested pairs and the classification, of which we report a compact excerpt in Table 1; and
  2. 2.
    up to six figures, one per case in Theorem 11, where we show in Figure 1.
In all cases, as Table 1 and Figure 1 clearly show, the prediction by Theorem 11 is accurate.
$\beta_{0}\beta_{1}>\lambda$
, $\beta_{1}\gg\beta_{0}$ (left).
$\beta_{0}\beta_{1}>\lambda$
, $|\beta_{0}-\beta_{1}|$ small (interior).
$\beta_{0}\beta_{1}>\lambda$
, $\beta_{0}\gg\beta_{1}$ (right).
$\beta_{0}\beta_{1}<\lambda$
, $\beta_{1}\gg\beta_{0}$ (left).
$\beta_{0}\beta_{1}<\lambda$
, $|\beta_{0}-\beta_{1}|$ small (either).
$\beta_{0}\beta_{1}<\lambda$
, $\beta_{0}\gg\beta_{1}$ (right).
Figure 1: Canonical examples for the six non-degenerate cases of Theorem 11.
Table 1: Excerpt (20 rows) from the batch verification on the grid of $(\beta_{0},\beta_{1})$.
$c$ $\kappa$ $\beta_{0}$ $\beta_{1}$ regime subcase predicted numeric comparison $\arg(\lambda_{\min})$
0.300 2.000 8.00 0.20 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 8.00 0.98 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 8.00 1.58 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 0.98 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 1.58 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 5.64 2.00 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 4.00 4.00 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 2.00 5.64 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 0.20 0.20 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 0.20 0.98 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 1.58 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 2.38 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.98 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 1.58 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 2.38 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 0.60 0.60 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 0.98 0.98 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 1.58 1.58 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 2.38 2.38 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
[…]
Acknowledgements
This work is partially supported by Grant No. SGS07/PŘF/2025, University of Ostrava.

### v2

@@
-Minimisation of Laplacian eigenvalue with indefinite weight under inhomogeneous Robin boundary condition
+One-dimensional optimisation of indefinite-weight principal eigenvalues under inhomogeneous Robin conditions and a Schrödinger-type perturbation
 Abstract
-This paper explores a certain Laplacian eigenvalue optimisation problem with indefinite weight under inhomogeneous Robin boundary condition.
-The minimum principal eigenvalue is fully determined in one dimension by formulating the problem as a shape optimisation one.
-The result is verified numerically using a shooting method.
-pacs:
-MSC Classification]35P15, 35J25
+We study the minimisation of the positive principal eigenvalue for an indefinite-weight problem under inhomogeneous Robin boundary conditions.
+The model is motivated by diffusive logistic equations in spatially heterogeneous environments, where the weight describes allocatable favourable resources and the Robin parameters measure boundary loss.
+After recalling the variational setting and the bang–bang reduction, we solve the one-dimensional optimisation problem completely: the optimal favourable set is an interval, and we determine whether it attaches to the left endpoint, to the right endpoint, or lies in the interior according to explicit criteria involving the two boundary parameters.
+The classification is then verified numerically by a transfer–matrix shooting method.
+We also introduce a Schrödinger-type extension with a fixed nonnegative background potential.
+In the coercive case we establish the corresponding principal-eigenvalue and bang–bang results, and in one dimension with constant potential we prove a stability result showing that minimisers for small background potential converge to those of the unperturbed problem.
+This perturbed regime is likewise illustrated numerically.
+Keywords: principal eigenvalue; indefinite weight; Robin boundary condition; shape optimisation; transfer–matrix method; Schrödinger-type perturbation
+MSC Classification: 35P15, 35J25, 49Q10, 92D25
 1 Introduction
-In this paper, we consider the following eigenvalue problem
-$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&\ \text{in}&\ \Omega\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&\ \text{on}&\ \partial\Omega\end{aligned}\right$. (1)
-where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ with Lipschitz boundary, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, and the weight function $m$ is a bounded measurable function that changes sign in $\Omega$, i.e.
-$\Omega_{m}^{+}:=\{x\in\Omega;\ m(x)>0\}$ satisfies $|\Omega_{m}^{+}|\in(0,|\Omega|)$, and that
-$1\leq m(x)\leq\kappa\ \text{a.e.}\ x\in\Omega$
-for a constant $\kappa>0$. The inhomogeneous Robin boundary condition applies, with $\beta:\partial\Omega\to\mathbb{R}$ a piecewise continuous function.
-According to [1], [2] and [3], $\lambda$ is called a principal eigenvalue of (1), if the associated eigenfunction $\phi$ is positive, where $\phi\in H^{1}(\Omega)$.
-The investigation of Laplacian eigenvalues with an indefinite weight under Robin boundary conditions originates from studies of diffusive logistics models in a bounded domain $\Omega$ with suitable boundary.
-In [4], it is shown that the density $u(x,t)$ of the species at location $x$ and time $t$ satisfies the following equation:
-$\begin{cases}u_{t}=\Delta u+\omega u[m(x)-u]\ \text{in}\ \Omega\times\mathbb{R}^{+},\\ \partial_{\mathbf{n}}u+\beta_{0}u=0\ \text{on}\ \partial\Omega\times\mathbb{R}^{+},\\ u(x,0)\geq 0,\ u(x,0)\not\equiv 0\ \text{in}\ \bar{\Omega},\end{cases}$ (2)
-where $\omega>0$ is a certain parameter.
-The Robin boundary parameter $\beta_{0}$ represents the inhospitableness of $\partial\Omega$.
-If $\beta_{0}=0$, the boundary is a barrier; if $\beta_{0}=+\infty$, the boundary is hostile.
-$m$ represents the intrinsic growth rate of the species, changing sign in $\Omega$, so that $\int_{\Omega}m$ represents the total resources in the spatially heterogeneous environment.
-In [5], the relation between the minimal positive principal eigenvalue $\lambda=\lambda(m)$ of (1) and the parameter $\omega$ in (2) is discovered: if $\omega\leq\lambda(m)$, then $u(x,t)\to 0$ uniformly in $\bar{\Omega}$ as $t\to\infty$ for proper initial data, i.e.
-the species extinguish; if $\omega>\lambda(m)$, then $u(x,t)\to u^{*}(x)$ uniformly in $\bar{\Omega}$ as $t\to\infty$, where $u^{*}$ is the unique positive steady solution of (2) in $W^{2,q}(\Omega)$ for every $q>1$, i.e.
-the species survives.
-Consequently, it is natural to consider the problem of minimising such an eigenvalue.
-In this paper we consider, more generally, the inhomogeneous Robin boundary condition, and we focus on the case $\beta\geq 0$ on $\partial\Omega$.
-We first prove the existence and uniqueness of a positive principal eigenvalue following the method by [6].
-Lemma 1 ([7], Theorem 11.4 and 11.10).
+Spatial heterogeneity, diffusion, and boundary exchange are basic features of many reaction–diffusion models arising in population dynamics, spatial ecology, and related applied fields [24, 21, 20, 7, 13].
+In such models, dispersal is described by diffusion, local growth or decay depends on the position $x\in\Omega$, and the interaction with the exterior is encoded through boundary conditions.
+A standard prototype is the diffusive logistic equation on a bounded habitat,
+$\begin{cases}\displaystyle u_{t}=\Delta u+\omega u[m(x)-u]&\text{in }\Omega\times\mathbb{R}_{+},\\[1.72218pt] \partial_{\mathbf{n}}u+\beta_{0}u=0&\text{on }\partial\Omega\times\mathbb{R}_{+},\\[1.72218pt] u(x,0)\geq 0,\quad u(x,0)\not\equiv 0&\text{in }\overline{\Omega},\end{cases}$ (1)
+where $\omega>0$ measures the overall reproductive intensity, $m$ is the intrinsic growth rate, and $\beta_{0}$ measures the inhospitableness of the boundary.
+When $\beta_{0}=0$ the boundary is impermeable, while positive values of $\beta_{0}$ model leakage, exchange, or mortality at the boundary; such Robin conditions are also standard in broader diffusion, thermal, and quantum models [8, 11, 19, 9].
+In a heterogeneous habitat, the sign-changing coefficient $m$ distinguishes favourable and unfavourable regions, and the principal eigenvalue of the linearised operator determines the threshold between extinction and persistence [6, 25, 3, 18].
+This leads to the indefinite-weight eigenvalue problem
+$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&&\text{in }\Omega,\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&&\text{on }\partial\Omega,\end{aligned}\right$. (2)
+where $\Omega\subset\mathbb{R}^{n}$ is a bounded Lipschitz domain, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, the weight $m\in L^{\infty}(\Omega)$ changes sign in $\Omega$, and
+$1\leq m(x)\leq\kappa\qquad\text{for a.e. }x\in\Omega,$
+for some fixed constant $\kappa>0$.
+Throughout the paper, $\beta:\partial\Omega\to\mathbb{R}$ is a piecewise continuous Robin parameter.
+Following [5, 23, 1, 12], a real number $\lambda$ is called a principal eigenvalue of (2) if the associated eigenfunction $\phi\in H^{1}(\Omega)$ is positive.
+The optimisation problem appears when the total amount of favourable habitat is limited. One imposes pointwise bounds on $m$ together with an integral constraint on the total resource,
+$\int_{\Omega}m\leq-m_{0}|\Omega|,$
+with $m_{0}$ fixed, and then asks how the favourable part of the habitat should be placed so as to minimise the positive principal eigenvalue.
+From the modelling point of view, the amount of favourable resource is prescribed, but its spatial arrangement is free; lowering the principal eigenvalue lowers the persistence threshold and therefore enlarges the range of viable reproduction parameters.
+Questions of this type belong to the broader spectral-optimisation tradition going back at least to Bôcher and Kreĭn [4, 16].
+In the indefinite-weight setting, the modern variational theory was developed for Neumann boundary conditions by Lou and Yanagida [18], for Robin boundary conditions by Hintermüller, Kao and Laurain [14], and further analysed by Lamboley, Laurain, Nadin and Privat [17].
+Beyond the interval case, complete structural results remain rare; the best understood higher-dimensional examples include cylindrical domains [15] and spherical shells [22].
+The first aim of the present paper is to treat the problem under inhomogeneous Robin boundary conditions in one dimension, allowing distinct nonnegative parameters at the two endpoints.
+This asymmetric setting is natural when the two ends of the habitat interact differently with the exterior.
+Our main theorem gives a complete classification of the optimal favourable interval: depending on the boundary parameters and the corresponding eigenvalue, the optimal interval may attach to the left endpoint, attach to the right endpoint, or lie in the interior.
+We then verify this classification numerically by a transfer–matrix shooting method.
+The second aim is to record a natural Schrödinger-type extension in which the optimisable indefinite weight is accompanied by a fixed hostile background potential.
+In ecological terms, this additional potential may represent mortality, harvesting, pollution, predation, or some other adverse effect that is not itself subject to optimisation [10].
+In the coercive case we show that the principal-eigenvalue theory and the bang–bang structure remain valid.
+We then study the one-dimensional family with constant potential and prove a stability result showing that minimisers for small positive potential converge to those of the unperturbed problem.
+A second numerical study, again based on transfer matrices, confirms this perturbative picture for representative canonical regimes.
+The paper is organised as follows.
+After the variational preliminaries and the bang–bang reduction in the remainder of this section, Section 2 contains the complete one-dimensional analysis for the original indefinite-weight problem.
+Section 3 introduces the Schrödinger-type extension, first at the general variational level and then for the one-dimensional constant-potential family.
+Section 4 presents the two numerical verification procedures: first for the unperturbed one-dimensional classification, and then for the perturbative constant-potential regime.
+We conclude with a brief discussion of the analytical scope of the results, their modelling interpretation, and possible directions for further work.
+We first prove the existence and uniqueness of a positive principal eigenvalue following the method in [1].
+Lemma 1.1 ([25], Theorem 11.4 and 11.10).
 For $\phi\in H^{1}(\Omega)$, let
@@
 Then $\lambda$ is a principal eigenvalue of (2), if and only if $\mu(\lambda)=0$.
-Lemma 2 ([6], Lemma 2).
+Lemma 1.2 ([1], Lemma 2).
 If $\phi_{0}\in H^{1}(\Omega)$ is a minimiser of $Q_{\phi}(\lambda)$, i.e. $\mu(\lambda)=\frac{Q_{\phi_{0}}(\lambda)}{\int_{\Omega}\phi_{0}^{2}}$, then
@@
 In other words, $\mu^{\prime\prime}(\lambda)\leq 0$.
-Meanwhile, it is obvious that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$.
+Meanwhile, it is clear that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$.
 Thus, $\mu=\mu(\lambda)$ increases until it attains its maximum, and decreases thereafter.
@@
 $Q_{\phi_{c}}(0)=0+c^{2}\int_{\partial\Omega}\beta=0,$
-and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 2, we get:
+and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 1.2, we get:
   • if $\int_{\Omega}m<0$, then $\mu^{\prime}(0)>0$, and $\mu$ has a unique positive zero corresponding to the unique positive principal eigenvalue of (2);
   • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)<0$, and $\mu$ has a unique negative zero corresponding to the unique negative principal eigenvalue of (2);
-  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (1) has a unique principal eigenvalue 0.
+  • if $\int_{\Omega}m=0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (2) has a unique principal eigenvalue 0.
 It is also well-known (cf. [5], [23], [18] and [17]) that the eigenvalue can be expressed with a Rayleigh quotient. To sum up, we have:
-Theorem 3.
+Theorem 1.3.
 If $\beta\geq 0$ on $\partial\Omega$ and $\beta>0$ on a subset of positive measure, the problem (2) admits a unique positive principal eigenvalue $\lambda$, depending on $m$. Moreover,
@@
 with respect to Robin parameter $\beta$.
-Theorem 4.
+Theorem 1.4.
 Suppose $\lambda_{1},\lambda_{2}$ are principal eigenvalues of problem (2) with Robin parameters $\beta_{1}$, $\beta_{2}$, respectively.
@@
 Proof.
-By Lemma 1,
+By Lemma 1.1,
 $\int_{\Omega}|\nabla\phi_{1}|^{2}+\int_{\partial\Omega}\beta_{1}\phi_{1}^{2}-\lambda_{1}\int_{\Omega}m\phi_{1}^{2}$
@@
 The following result is proven in [18] for Neumann boundary condition ($\beta\equiv 0$), but the extension to general Robin case is straightforward:
-Theorem 5.
+Theorem 1.5.
 There exists some weight function $m^{*}\in\mathcal{M}_{m_{0},\kappa}$ such that the infimum of (3) is attained at $\lambda(m^{*})$.
@@
 2 The 1-dimensional case
-In this section, we derive the complete solution of the problem (1) in one dimension.
+In this section, we solve problem (2) completely in one dimension.
 Consider the problem (2) in $\Omega=(0,1)\subseteq\mathbb{R}$. The following result is shown in [17] for homogeneous Robin conditions, but its generalisation is straightforward.
-Proposition 6.
+Proposition 2.1.
 If $\Omega=(0,1)$, then any optimal set $E^{*}$ for the problem (5) is an interval.
 Now let the optimal set be the interval $E=(a,b)$ where $0<a<b<1$ and c=bac=b-a.
-We intend to find the optimal value of $a$ following the method by [2].
+We determine the optimal value of $a$ by adapting the method of [14].
 Let $\beta_{0}=\beta(0)\geq 0$ and $\beta_{1}=\beta(1)>0$.
@@
 (7)
-where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. Now we simplify the boundary and transmission conditions in (6). By organising the coefficients of $C_{i}$ ($1\leq i\leq 6$), we get
+where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. We now simplify the boundary and transmission conditions in (6). Collecting the coefficients of $C_{i}$ ($1\leq i\leq 6$), we obtain
 $\left\{\begin{aligned} &\left(-\sqrt{\lambda}\sinh(a\sqrt{\lambda})-\beta_{0}\cosh(a\sqrt{\lambda})\right)C_{1}+\left(\sqrt{\lambda}\cosh(a\sqrt{\lambda})+\beta_{0}\sinh(a\sqrt{\lambda})\right)C_{2}=0\\ &\left(\sqrt{\lambda}\sinh(\sqrt{\lambda}(1-b))+\beta_{1}\cosh(\sqrt{\lambda}(1-b))\right)C_{5}\\ &\ \ \ +\left(\sqrt{\lambda}\cosh(\sqrt{\lambda}(1-b))+\beta_{1}\sinh(\sqrt{\lambda}(1-b))\right)C_{6}=0\\ &C_{3}-C_{1}=0\\ &C_{5}-C_{3}\cos(\sqrt{\lambda\kappa}c)-C_{4}\sin(\sqrt{\lambda\kappa}c)=0\\ &\sqrt{\kappa}C_{4}-C_{2}=0\\ &C_{6}+C_{3}\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)-C_{4}\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)=0\end{aligned}\right$.
@@
 $\left(-\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(a)-\beta_{0}{\mathfrak{c}}_{\lambda}(a)\right)\left[\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ +\left.\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]\\ =\sqrt{\kappa}\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(a)+\beta_{0}\sinh(a\sqrt{\lambda})\right)\left[\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ -\left.\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right].$ (10)
-First we order the term and get
+Rearranging the terms gives
 $\cosh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-(\beta_{0}+\beta_{1})\sqrt{\lambda\kappa}\cos(\sqrt{\lambda\kappa}c)+(\lambda\kappa-\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{0}-\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right]\\ +\sinh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{1}-\beta_{0})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\beta_{0}\beta_{1}\kappa-\lambda)\sin(\sqrt{\lambda\kappa}c)\right)\right]=0.$ (11)
-Let us first verify (11) in two extreme cases.
-When $\beta_{0}=\beta_{1}=0$, the problem is reduced to Neumann boundary condition, and we have obtained the characteristic equation [^2] We remark that the characteristic equations (18) and (20) obtained in [2] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
+It is useful to check (11) in two extreme cases.
+When $\beta_{0}=\beta_{1}=0$, the problem reduces to the Neumann case, and we obtain the characteristic equation [^5] We remark that the characteristic equations (18) and (20) obtained in [14] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
 $\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-\tanh(\lambda^{\frac{1}{2}}a)}{1+\kappa^{-\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)},$ (12)
@@
 $\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (13)
-On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem is reduced to Dirichlet boundary condition, we similarly get the characteristic equation
+On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem reduces to the Dirichlet case, and we similarly obtain the characteristic equation
 $\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{-\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)+\tanh(\lambda^{\frac{1}{2}}a)}{\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-1},$ (14)
@@
 $\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=-\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (15)
-Now we use the product-to-sum formulae for hyperbolic functions to simplify it. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
+We next use the product-to-sum formulae for hyperbolic functions to simplify this expression. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
 $0=(\kappa+1)(\lambda-\beta_{0}\beta_{1})\cosh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +(\kappa+1)(\beta_{0}-\beta_{1})\sqrt{\lambda}\sinh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +\cosh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\lambda+\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)\right)\\ +\sinh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\beta_{0}+\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\kappa}(\beta_{0}\beta_{1}+\lambda)\cos(\sqrt{\lambda\kappa}c)\right).$ (16)
@@
 $g(a,\beta_{0},\beta_{1},\lambda):=(\lambda-\beta_{0}\beta_{1})\tanh\left(2\sqrt{\lambda}\left(a-\frac{1-c}{2}\right)\right)+\sqrt{\lambda}(\beta_{0}-\beta_{1}).$ (19)
-Lemma 7.
+Lemma 2.2.
 For any $a\in[0,1-c]$, $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\not\equiv 0$, where $\kappa>0$.
@@
 and comparing (22) and (25) we get $\kappa=0$, a contradiction to the assumption that $\kappa>0$. ∎
-Lemma 8.
+Lemma 2.3.
 Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where
@@
 We now prove:
-Proposition 9.
+Proposition 2.4.
 Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is given in (26).
@@
 $\partial_{\lambda}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)=-\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda).$ (33)
-It is shown in Theorem 4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
+It is shown in Theorem 1.4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
 $\partial_{\beta_{1}}f>0\Rightarrow\partial_{\lambda}f<0\Rightarrow\partial_{a}f\cdot\partial_{a}\lambda>0\Leftrightarrow g\cdot\partial_{a}\lambda>0,$ (34)
@@
 $A(a)=-2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(2a+c-1)\big),\\[5.69054pt] B(a)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(2a+c-1)\big).$ (37)
-We prove in Lemma 7 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
+We prove in Lemma 2.2 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
 $\beta_{0}=\beta_{0}^{*}(a):=-\frac{B(a)}{A(a)}.$ (38)
@@
 Denote by $h=h(c,\kappa,\lambda)$ the expression on the right-hand side of (39).
-We show in Lemma 8 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$.
+We show in Lemma 2.3 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$.
 In other words, $\partial_{\beta_{1}}f<0$ if $\beta_{0}>\beta_{0}^{*}$ and $\partial_{\beta_{1}}f>0$ if $\beta_{0}<\beta_{0}^{*}$.
 ∎
-Lemma 10.
+Lemma 2.5.
 Denote by
@@
 $A(a)\geq 2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ +(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\\ \geq-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)>0,$ (41)
-where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 8. Meanwhile,
+where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 2.3. Meanwhile,
 $B(a)\leq B(0)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ \leq 2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)$ (42)
@@
 $\beta_{0}^{*}\leq\frac{\sqrt{\kappa}\pi}{c}\frac{\sinh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)}{\kappa+1-(\kappa-1)\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)},$ (45)
-where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 8. ∎
-Now we are ready to present the complete result on the 1-dimensional interval.
-Theorem 11.
+where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 2.3. ∎
+We are now in a position to state the complete result for the one-dimensional interval.
+Theorem 2.6.
 Consider problem (5) in $\Omega=(0,1)\subseteq\mathbb{R}^{1}$ with $\beta_{0}>\tilde{\beta_{0}^{*}}$, and $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is as in (26) and $\tilde{\beta_{0}^{*}}$ is as in (40).
@@
 Now we consider the case $\beta_{0}\beta_{1}>\lambda$.
-From (19), it is obvious that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46).
+From (19), it is clear that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46).
 We now need to check whether $a^{*}\in[0,1-c]$.
@@
 Similarly we consider the case $\beta_{0}\beta_{1}<\lambda$.
-From (19), it is obvious that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46).
+From (19), it is clear that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46).
 We now need to check whether $a^{*}\in[0,1-c]$.
@@
 then $g<0$ for $a\in(0,a^{*})$ and $g>0$ for $a\in(a^{*},1-c)$.
-This completes the proof together with Proposition 9. ∎
-Remark on homogeneous Robin boundary
+This completes the proof together with Proposition 2.4. ∎
+Remark on homogeneous Robin boundary.
 In [14], when studying the same problem under homogeneous Robin boundary ($\beta_{0}=\beta_{1}=\beta$), it is possible to reduce the condition $g=0$ to $\lambda=\beta_{0}\beta_{1}$, which could be plugged into $f=0$ to eliminate the terms involving $a$, which yields an equation of $\beta$, whose smallest positive solution can be expressed explicitly as $\beta^{*}=\beta^{*}(c,\kappa)$.
@@
 (19)) and (17).
-Nevertheless, in the next section we shall see that the criteria in Theorem 11 involving $\lambda$ could still be tested and thus the theorem itself be verified numerically.
-3 Transfer–matrix shooting
-In this section, we use a classical shooting method to numerically verify Theorem 11 for $\beta_{0}\beta_{1}\neq\lambda$.
+Nevertheless, in Section 4 we shall see that the criteria in Theorem 2.6 involving $\lambda$ can still be tested, and thus the theorem itself can be verified numerically.
+3 A Schrödinger-type extension with a fixed background potential
+3.1 General variational and bang–bang results
+The ecological interpretation of (2) suggests a natural extension in which the sign-changing weight $m$ is still the optimisable resource, but an additional spatially varying hostile field is prescribed in advance.
+More precisely, let $q\in L^{\infty}(\Omega)$ and consider
+$\left\{\begin{aligned} \Delta\phi+(\lambda m(x)-q(x))\phi&=0&&\text{in }\Omega,\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&&\text{on }\partial\Omega.\end{aligned}\right$. (53)
+When $q\equiv 0$, this reduces to (2).
+We think of $q$ as a fixed background penalty: depending on the application, it may represent mortality, harvesting pressure, pollution, predation, or some other adverse environmental effect that is not itself subject to optimisation.
+In terms of the diffusive logistic model, the corresponding parabolic equation is
+$\begin{cases}\displaystyle u_{t}=\Delta u+\bigl(\omega m(x)-q(x)\bigr)u-u^{2}&\text{in }\Omega\times\mathbb{R}_{+},\\[1.72218pt] \partial_{\mathbf{n}}u+\beta_{0}u=0&\text{on }\partial\Omega\times\mathbb{R}_{+},\end{cases}$ (54)
+whose linearisation at $u\equiv 0$ is exactly (53).
+Thus the optimisation problem asks how a favourable region should be placed when the habitat is already subject to a non-removable hostile background.
+This fits naturally with the reaction–diffusion literature on heterogeneous environments, boundary loss, and harvesting or protection effects; see, for example, [7, 8, 10].
+From the spectral point of view, (53) is a Schrödinger-type problem: equivalently, one may rewrite it as $\Delta\phi+q\phi=\lambda m\phi$ with Robin boundary condition.
+For related indefinite Robin eigenvalue problems on general domains, see also [12].
+In this section we record the part of the theory that survives with essentially the same proofs as in the case $q\equiv 0$. We restrict ourselves to the coercive situation
+$q\geq 0\quad\text{a.e. in }\Omega,\qquad\beta\geq 0\quad\text{on }\partial\Omega,$ (55)
+and assume that either $q>0$ on a subset of $\Omega$ of positive measure or $\beta>0$ on a subset of $\partial\Omega$ of positive surface measure.
+This avoids the degenerate pure Neumann case $(q,\beta)\equiv(0,0)$, where the sign of $\int_{\Omega}m$ again becomes decisive.
+For $\phi\in H^{1}(\Omega)$ and $\lambda\in\mathbb{R}$, set
+$Q^{q}_{\phi}(\lambda):=\int_{\Omega}|\nabla\phi|^{2}+\int_{\Omega}q\phi^{2}+\int_{\partial\Omega}\beta\phi^{2}-\lambda\int_{\Omega}m\phi^{2},$
+and
+$\mu_{q}(\lambda):=\inf_{\phi\in H^{1}(\Omega)}\frac{Q^{q}_{\phi}(\lambda)}{\int_{\Omega}\phi^{2}}.$
+Exactly as in Lemma 1.1 and Lemma 1.2, one shows that $\lambda$ is a principal eigenvalue of (53) if and only if $\mu_{q}(\lambda)=0$, and that if $\phi_{0}$ is a minimiser in the definition of $\mu_{q}(\lambda)$, then
+$\mu_{q}^{\prime}(\lambda)=-\frac{\int_{\Omega}m\phi_{0}^{2}}{\int_{\Omega}\phi_{0}^{2}}.$
+Since these arguments are literally the same as before, we omit the proof.
+Theorem 3.1.
+Assume (55), and assume further that either $q>0$ on a subset of $\Omega$ of positive measure or $\beta>0$ on a subset of $\partial\Omega$ of positive surface measure.
+Let $m\in L^{\infty}(\Omega)$ be sign-changing with
+$1\leq m(x)\leq\kappa\qquad\text{for a.e. }x\in\Omega.$
+Then problem (53) admits a unique positive principal eigenvalue, denoted by $\lambda_{q}(m)$. Moreover,
+$\lambda_{q}(m)=\inf_{\phi\in S(m)}\frac{\displaystyle\int_{\Omega}|\nabla\phi|^{2}+\int_{\Omega}q\phi^{2}+\int_{\partial\Omega}\beta\phi^{2}}{\displaystyle\int_{\Omega}m\phi^{2}},\qquad S(m):=\left\{\phi\in H^{1}(\Omega);\ \int_{\Omega}m\phi^{2}>0\right\}.$ (56)
+Proof.
+For every fixed $\phi\in H^{1}(\Omega)$, the map $\lambda\mapsto Q^{q}_{\phi}(\lambda)$ is affine, hence concave.
+Therefore $\mu_{q}$ is concave as the infimum of affine functions.
+Since $m$ changes sign, there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ such that
+$\int_{\Omega}m\phi_{1}^{2}>0,\qquad\int_{\Omega}m\phi_{2}^{2}<0,$
+and consequently $\mu_{q}(\lambda)\to$-$\infty$ both as $\lambda\to+\infty$ and as $\lambda\to$-$\infty$.
+Next consider
+$\Lambda_{1}(q,\beta):=\inf_{\phi\in H^{1}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\Omega}|\nabla\phi|^{2}+\int_{\Omega}q\phi^{2}+\int_{\partial\Omega}\beta\phi^{2}}{\displaystyle\int_{\Omega}\phi^{2}}.$
+By the compact embedding $H^{1}(\Omega)\hookrightarrow L^{2}(\Omega)$, the infimum is attained.
+Since $q\geq 0$ and $\beta\geq 0$, the numerator is nonnegative.
+If $\Lambda_{1}(q,\beta)=0$, then for a minimiser $\phi$ we must have
+$\nabla\phi=0\quad\text{a.e. in }\Omega,\qquad q\phi=0\quad\text{a.e. in }\Omega,\qquad\beta\phi=0\quad\text{a.e. on }\partial\Omega.$
+Hence $\phi$ is constant, and the additional assumption that either $q>0$ or $\beta>0$ on a set of positive measure forces this constant to be zero, a contradiction.
+Therefore $\Lambda_{1}(q,\beta)>0$, and so
+$\mu_{q}(0)\geq\Lambda_{1}(q,\beta)>0.$
+By concavity, $\mu_{q}$ has exactly two zeros, one positive and one negative. The positive zero is the unique positive principal eigenvalue of (53), which we denote by $\lambda_{q}(m)$.
+Finally, since $\mu_{q}\bigl(\lambda_{q}(m)\bigr)=0$, the same argument as in Theorem 1.3 yields the Rayleigh quotient formula (56). ∎
+We now consider the optimisation problem
+$\inf_{m\in\mathcal{M}_{m_{0},\kappa}}\lambda_{q}(m),\qquad\mathcal{M}_{m_{0},\kappa}:=\left\{m\in L^{\infty}(\Omega);\ -1\leq m\leq\kappa,\ |\Omega_{m}^{+}|>0,\ \int_{\Omega}m\leq-m_{0}|\Omega|\right\},$ (57)
+where $m_{0}\in(-\kappa,1)$.
+Theorem 3.2.
+Under the assumptions of Theorem 3.1, the optimisation problem (57) admits a minimiser. Moreover, there exists a minimiser $m^{*}\in\mathcal{M}_{m_{0},\kappa}$ of the form
+$m^{*}=\kappa\chi_{E^{*}}-\chi_{\Omega\setminus E^{*}}\qquad\text{a.e. in }\Omega,$ (58)
+for some measurable set $E^{*}\subset\Omega$. In addition, the volume constraint is active,
+$\int_{\Omega}m^{*}=-m_{0}|\Omega|.$ (59)
+Equivalently, if
+$c:=\frac{1-m_{0}}{1+\kappa},$
+then the optimisation problem can be rewritten as
+$\inf_{E\subset\Omega,\ |E|=c|\Omega|}\lambda_{q}(E),\qquad\lambda_{q}(E):=\lambda_{q}\bigl(\kappa\chi_{E}-\chi_{\Omega\setminus E}\bigr).$ (60)
+Proof.
+The existence part follows by the same direct-method argument as in [18, Theorem 1.4] and [14, Theorem 1.2].
+Indeed, compared with the case $q\equiv 0$, the only modification is the additional fixed term $\int_{\Omega}q\phi^{2}$ in the numerator of the Rayleigh quotient, which is independent of the optimisation variable $m$ and does not affect the compactness or lower-semicontinuity arguments.
+Let $m^{*}$ be a minimiser, and let $\phi^{*}>0$ be an associated eigenfunction.
+By homogeneity we may normalise $\phi^{*}$ arbitrarily.
+Since $m^{*}$ minimises $\lambda_{q}$, for every $m\in\mathcal{M}_{m_{0},\kappa}$ we have
+$\lambda_{q}(m^{*})\leq\lambda_{q}(m)\leq\frac{\displaystyle\int_{\Omega}|\nabla\phi^{*}|^{2}+\int_{\Omega}q(\phi^{*})^{2}+\int_{\partial\Omega}\beta(\phi^{*})^{2}}{\displaystyle\int_{\Omega}m(\phi^{*})^{2}}.$
+Since equality holds when $m=m^{*}$, the minimiser $m^{*}$ maximises the linear functional
+$L_{\phi^{*}}(m):=\int_{\Omega}m(\phi^{*})^{2}$
+over $\mathcal{M}_{m_{0},\kappa}$.
+We first prove that the integral constraint is active. If
+$\int_{\Omega}m^{*}<-m_{0}|\Omega|,$
+then, because $\phi^{*}>0$ a.e.
+in $\Omega$ and $m^{*}\leq\kappa$, one can increase $m^{*}$ by a small positive amount on a subset of positive measure where $m^{*}<\kappa$, while staying inside $\mathcal{M}_{m_{0},\kappa}$.
+This strictly increases $L_{\phi^{*}}(m)$, contradicting the maximality of $m^{*}$.
+Hence (59) holds.
+Now consider the convex set
+$\widetilde{\mathcal{M}}_{m_{0},\kappa}:=\left\{m\in L^{\infty}(\Omega);\ -1\leq m\leq\kappa,\ \int_{\Omega}m=-m_{0}|\Omega|\right\}.$
+The functional $L_{\phi^{*}}$ is linear, hence it attains its maximum on an extreme point of $\widetilde{\mathcal{M}}_{m_{0},\kappa}$.
+The extreme points are exactly the functions of the form $\kappa\chi_{E}-\chi_{\Omega\setminus E}$ with
+$|E|=\frac{1-m_{0}}{1+\kappa}|\Omega|=c|\Omega|.$
+Therefore there exists a maximiser of $L_{\phi^{*}}$, and hence a minimiser of (57), of the bang-bang form (58). The identity (59) then implies $|E^{*}|=c|\Omega|$, which yields (60). ∎
+3.2 The one-dimensional interval family with constant background potential
+We now specialise (53) to $\Omega=(0,1)$ and to a constant background potential $q\equiv q_{0}\geq 0$.
+We do not attempt a full analogue of Theorem 2.6; instead, we record a perturbative result which already gives a rigorous bridge between the unperturbed problem and the Schrödinger-type model, and which is directly suited for numerical verification.
+Fix $c\in(0,1)$, let
+$E_{a}:=(a,a+c),\qquad a\in[0,1-c],$
+and define
+$m_{a}:=\kappa\chi_{E_{a}}-\chi_{(0,1)\setminus E_{a}}.$
+For every $a\in[0,1-c]$ and every $q_{0}\geq 0$, consider
+$\left\{\begin{aligned} \phi^{\prime\prime}+(\lambda m_{a}(x)-q_{0})\phi&=0&&\text{in }(0,1),\\ \phi^{\prime}(0)-\beta_{0}\phi(0)&=0,\\ \phi^{\prime}(1)+\beta_{1}\phi(1)&=0,\end{aligned}\right$. (61)
+where $\beta_{0}\geq 0$ and $\beta_{1}>0$. By Theorem 3.1, for every such pair $(a,q_{0})$ the problem (61) admits a unique positive principal eigenvalue; we denote it by $\lambda_{q_{0}}(a)$.
+The key observation is that the transfer–matrix formulation used in Section 4 extends with essentially no change in strategy.
+Proposition 3.3.
+For $m\in\{-1,\kappa\}$, set
+$Q_{m}^{(q_{0})}(\lambda):=\begin{pmatrix}0&1\\ q_{0}-\lambda m&0\end{pmatrix}$.
+Then (61) is equivalent to the characteristic equation
+$R_{q_{0}}(a,\lambda):=e_{2}M_{q_{0}}(a,\lambda)(1,\beta_{0})^{\top}+\beta_{1}e_{1}M_{q_{0}}(a,\lambda)(1,\beta_{0})^{\top}=0,$ (62)
+where
+$M_{q_{0}}(a,\lambda):=e^{(1-a-c)Q_{-1}^{(q_{0})}(\lambda)}e^{cQ_{\kappa}^{(q_{0})}(\lambda)}e^{aQ_{-1}^{(q_{0})}(\lambda)}.$ (63)
+Moreover, $\lambda_{q_{0}}(a)$ is the smallest positive root of (62).
+Proof.
+Let $w=(\phi,\phi^{\prime})^{\top}$. On every subinterval where $m_{a}$ is constant, the equation in (61) becomes
+$w^{\prime}=Q_{m}^{(q_{0})}(\lambda)w,$
+and therefore the solution is propagated by matrix exponentials.
+Since (61) is homogeneous in $\phi$, we may normalise $\phi(0)=1$, so that the left Robin condition gives $w(0)=(1,\beta_{0})^{\top}$.
+Propagating first across $(0,a)$, then $(a,a+c)$, and finally $(a+c,1)$ yields (63).
+The right Robin condition is exactly (62).
+The last assertion follows because the principal eigenvalue is the first eigenvalue of this regular one-dimensional problem.
+∎
+The next result is the basic perturbative statement needed for the Schrödinger extension.
+Proposition 3.4.
+Fix $\overline{q}>0$. Then the map
+$(a,q_{0})\longmapsto\lambda_{q_{0}}(a)$
+is continuous on $0,1-c]\times[0,\overline{q}]$. In particular, for every $q_{0}\in[0,\overline{q}]$ the minimiser set
+$\mathcal{A}(q_{0}):=\operatorname*{argmin}_{a\in[0,1-c]}\lambda_{q_{0}}(a)$
+is nonempty and compact.
+Proof.
+By Proposition 3.3, the characteristic function $R_{q_{0}}(a,\lambda)$ is real-analytic in $(a,q_{0},\lambda)$.
+For fixed $(a,q_{0})$, the principal eigenvalue of the regular boundary-value problem (61) is simple: indeed, if $\phi_{1}$ and $\phi_{2}$ correspond to the same eigenvalue, then their Wronskian is constant, and the left Robin condition gives $W(0)=\phi_{1}(0)\phi_{2}^{\prime}(0)-\phi_{1}^{\prime}(0)\phi_{2}(0)=0$, hence $W\equiv 0$ and $\phi_{1},\phi_{2}$ are linearly dependent.
+For separated one-dimensional boundary conditions this implies that $\lambda=\lambda_{q_{0}}(a)$ is a simple zero of $R_{q_{0}}(a,\lambda)$.
+The implicit function theorem therefore yields local continuous dependence of $\lambda_{q_{0}}(a)$ on $(a,q_{0})$, and hence continuity on the whole compact rectangle $0,1-c]\times[0,\overline{q}]$.
+The nonemptiness and compactness of $\mathcal{A}(q_{0})$ then follow from the continuity of $a\mapsto\lambda_{q_{0}}(a)$ on the compact interval [0,1c][0,1-c]. ∎
+Theorem 3.5.
+Let $q_{j}\to 0^{+}$, and for each $j$ choose some $a_{j}\in\mathcal{A}(q_{j})$.
+Then every accumulation point of $(a_{j})$ belongs to $\mathcal{A}(0)$.
+Equivalently, every family of minimisers for the constant-potential problem converges, along subsequences, to minimisers of the unperturbed problem.
+Proof.
+Since $a_{j}\in[0,1-c]$, the sequence $(a_{j})$ has an accumulation point; passing to a subsequence if necessary, we may assume that $a_{j}\to a_{*}\in[0,1-c]$.
+Let $a\in[0,1-c]$ be arbitrary.
+Since each $a_{j}$ minimises $\lambda_{q_{j}}$, we have
+$\lambda_{q_{j}}(a_{j})\leq\lambda_{q_{j}}(a).$
+Now let $j\to\infty$. By Proposition 3.4,
+$\lambda_{0}(a_{*})=\lim_{j\to\infty}\lambda_{q_{j}}(a_{j})\leq\lim_{j\to\infty}\lambda_{q_{j}}(a)=\lambda_{0}(a).$
+Since $a\in[0,1-c]$ was arbitrary, this shows that $a_{*}\in\mathcal{A}(0)$. ∎
+Combining Theorem 3.5 with the one-dimensional classification from Section 2, we immediately obtain the following consequence.
+Corollary 3.6.
+Assume that the parameters satisfy the hypotheses of Theorem 2.6, and let $q_{j}\to 0^{+}$. For each $j$, choose $a_{j}\in\mathcal{A}(q_{j})$.
+  1. (i)
+    If Theorem 2.6 yields the unique minimiser $a=0$, then $a_{j}\to 0$.
+  2. (ii)
+    If Theorem 2.6 yields the unique minimiser a=1ca=1-c, then $a_{j}\to 1-c$.
+  3. (iii)
+    If Theorem 2.6 yields the unique interior minimiser $a=a^{*}$, then $a_{j}\to a^{*}$.
+  4. (iv)
+    If Theorem 2.6 yields exactly the two minimisers $a=0$ and a=1ca=1-c, then
+    $\operatorname{dist}\bigl(a_{j},\{0,1-c\}\bigr)\to 0.$
+In the fully degenerate case where every $a\in[0,1-c]$ is optimal for the unperturbed problem, Theorem 3.5 gives no further restriction, as expected.
+Proof.
+By Theorem 3.5, every accumulation point of $(a_{j})$ is a minimiser of the unperturbed problem.
+When Theorem 2.6 gives a unique minimiser, the whole sequence must converge to it.
+When the unperturbed minimiser set is $\{0,1-c\}$, every accumulation point belongs to this two-point set, which is equivalent to the distance statement.
+∎
+Corollary 3.6 is the precise sense in which the one-dimensional classification obtained for $q_{0}=0$ persists under a small constant Schrödinger perturbation.
+It also gives a clean target for the numerical part: one should observe the minimisers of $a\mapsto\lambda_{q_{0}}(a)$ drifting continuously towards the locations predicted by Theorem 2.6 as $q_{0}\downarrow 0$.
+4 Numerical verification by a transfer–matrix shooting method
+This section collects the numerical checks accompanying the two one-dimensional parts of the paper.
+We first verify Theorem 2.6 for the unperturbed problem, and then we verify Corollary 3.6 for the constant-potential perturbation.
+In both cases the computation is based on the same transfer–matrix shooting principle.
+The code and machine-readable outputs used to generate the numerical results in this section are available in [26].
+4.1 Numerical verification of Theorem 2.6
+In this subsection, we use a classical shooting method to numerically verify Theorem 2.6 for $\beta_{0}\beta_{1}\neq\lambda$.
 For given $(c,\kappa,\beta_{0},\beta_{1})$ and a placement $a\in[0,1-c]$ of the favourable subinterval $(a,a+c)$, the equation
@@
 Finally, such numerical computation is done for a finite grid of $a$-values in the interval [0,1c][0,1-c], with a chosen reasonable resolution $N_{a}$, in order to produce the curve $a\mapsto\lambda(a)$ and to determine where the eigenvalue is minimised.
-To compare the numerical result with Theorem 11 (non-degenerate cases $\beta_{0}\beta_{1}\neq\lambda$ only), we run the computation for a range of values for $(\beta_{0},\beta_{1})$ and for each pair of the Robin parameters, we verify the criteria in Theorem 11 for all computed eigenvalues $\lambda(a_{j})$, where $\{a_{j}\}_{j}$ are the grid of $a$-values in the interval [0,1c][0,1-c].
-If a certain criterion in Theorem 11 is satisfied by all such $\lambda(a_{j})$, we classify the pair $(\beta_{0},\beta_{1})$ in the case corresponding to the valid criterion in Theorem 11, otherwise we do not classify this pair.
-For all those pairs of Robin parameters that are classified in a certain case of Theorem 11, we compare the predicted location for the minimum eigenvalue and the numerical result, and thus verify the correctness of Theorem 11.
+To compare the numerical result with Theorem 2.6 (non-degenerate cases $\beta_{0}\beta_{1}\neq\lambda$ only), we run the computation for a range of values for $(\beta_{0},\beta_{1})$ and for each pair of the Robin parameters, we verify the criteria in Theorem 2.6 for all computed eigenvalues $\lambda(a_{j})$, where $\{a_{j}\}_{j}$ are the grid of $a$-values in the interval [0,1c][0,1-c].
+If a certain criterion in Theorem 2.6 is satisfied by all such $\lambda(a_{j})$, we classify the pair $(\beta_{0},\beta_{1})$ in the case corresponding to the valid criterion in Theorem 2.6, otherwise we do not classify this pair.
+For all those pairs of Robin parameters that are classified in a certain case of Theorem 2.6, we compare the predicted location for the minimum eigenvalue and the numerical result, and thus verify the correctness of Theorem 2.6.
 Choice of parameters
@@
 The upper bound $\beta_{\max}$ is chosen to ensure that $\beta_{0}\beta_{1}-\lambda$ changes sign for the chosen $\kappa=2.0$ and $c=0.3$.
-In general, if some case from Theorem 11 is not found in a given $\beta$–grid, it is appropriate to widen $\beta_{\min},\beta_{\max}]$ or increase $n_{\beta}$.
+In general, if some case from Theorem 2.6 is not found in a given $\beta$–grid, it is appropriate to widen $\beta_{\min},\beta_{\max}]$ or increase $n_{\beta}$.
 Outputs
@@
   2. 2.
-    up to six figures, one per case in Theorem 11, where we show in Figure 1.
-In all cases, as Table 1 and Figure 1 clearly show, the prediction by Theorem 11 is accurate.
+    up to six figures, one per case in Theorem 2.6, where we show in Figure 1.
+In all cases, as Table 1 and Figure 1 clearly show, the prediction by Theorem 2.6 is accurate.
 $\beta_{0}\beta_{1}>\lambda$
@@
 , $\beta_{0}\gg\beta_{1}$ (right).
-Figure 1: Canonical examples for the six non-degenerate cases of Theorem 11.
+Figure 1: Canonical examples for the six non-degenerate cases of Theorem 2.6.
 Table 1: Excerpt (20 rows) from the batch verification on the grid of $(\beta_{0},\beta_{1})$.
@@
 0.300 2.000 2.38 2.38 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
-[…]
-Acknowledgements
-This work is partially supported by Grant No. SGS07/PŘF/2025, University of Ostrava.
+4.2 Numerical verification of Corollary 3.6
+We now verify Corollary 3.6 by the same transfer–matrix strategy as above. For fixed $q_{0}\geq 0$ and $a\in[0,1-c]$, Proposition 3.3 reduces the eigenvalue problem (61) to the scalar equation
+$R_{q_{0}}(a,\lambda)=0,$
+and the principal eigenvalue $\lambda_{q_{0}}(a)$ is computed as the smallest positive root.
+Numerically, this is done by a uniform scan in $\lambda$, followed by bisection on the leftmost interval where the residual changes sign.
+The only modification with respect to Section 4 is that the matrices $Q_{m}$ are replaced by
+$Q_{m}^{(q_{0})}(\lambda)=\begin{pmatrix}0&1\\ q_{0}-\lambda m&0\end{pmatrix}$.
+For the placement variable, we work on a uniform grid $\{a_{j}\}_{j=0}^{N_{a}}\subset[0,1-c]$ and define the discrete minimiser set
+$\mathcal{A}_{N_{a}}(q_{0}):=\arg\min\{\lambda_{q_{0}}(a_{j}):0\leq j\leq N_{a}\}.$
+To compare with Corollary 3.6, we also record the deterministic selector
+$a^{\sharp}(q_{0}):=\min\mathcal{A}_{N_{a}}(q_{0}),$
+that is, the leftmost point in the discrete argmin set.
+If the unperturbed minimiser is unique, then $a^{\sharp}(q_{0})$ should approach it as $q_{0}\downarrow 0$; in the two-endpoint case, the distance from $a^{\sharp}(q_{0})$ to $\{0,1-c\}$ should tend to 0.
+Choice of parameters.
+We take $c=0.3$, $\kappa=2.0$, $N_{a}=240$ subintervals in the $a$-grid (so $241$ sampled $a$-values), and an adaptive scan in $\lambda$ starting from $500$ sample points before bisection.
+The values of the background potential are
+$q_{0}\in\{0.2,0.1,0.05,0.02,0.01,0\}.$
+We test the following four canonical pairs $(\beta_{0},\beta_{1})$ corresponding to the four alternatives in Corollary 3.6:
+$(0.2,2.8),\qquad(2.8,0.2),\qquad(3.67,3.67),\qquad(0.2,0.2).$
+These respectively represent a unique left endpoint minimiser, a unique right endpoint minimiser, a unique interior minimiser, and the two-endpoint minimiser set.
+Table 2: Numerical verification of Corollary 3.6 for four canonical parameter choices.
+In each row, the displayed discrete argmin set is the same for every sampled value $q_{0}\in\{0.2,0.1,0.05,0.02,0.01,0\}$.
+case $(\beta_{0},\beta_{1})$ predicted $\mathcal{A}(0)$ observed $\mathcal{A}_{N_{a}}(q_{0})$ $\max\operatorname{dist}(a^{\sharp}(q_{0}),\mathcal{A}(0))$
+unique left endpoint $(0.2,2.8)$ $\{0\}$ always $\{0\}$ 0
+unique right endpoint $(2.8,0.2)$ $\{1-c\}=\{0.7\}$ always $\{0.7\}$ 0
+unique interior point $(3.67,3.67)$ $\{\tfrac{1-c}{2}\}=\{0.35\}$ always $\{0.35\}$ 0
+two endpoint minimisers $(0.2,0.2)$ $\{0,1-c\}=\{0,0.7\}$ always $\{0,0.7\}$ 0
+Table 2 already confirms the conclusion of Corollary 3.6 in these four representative situations.
+In fact, for the chosen samples, the discrete minimiser set does not merely converge to the unperturbed one as $q_{0}\downarrow 0$; it coincides with the prediction for every tested value of $q_{0}$.
+The endpoint pair $(\beta_{0},\beta_{1})=(0.2,0.2)$ is especially transparent: for each sampled $q_{0}$, the two boundary placements remain tied, so the discrete argmin set is exactly $\{0,1-c\}$.
+The corresponding eigenvalue profiles are shown in Figure 2.
+For each canonical choice, we plot $a\mapsto\lambda_{q_{0}}(a)$ for $q_{0}=0.2$, $0.05$, and 0.
+The figure makes the four qualitative behaviours visually clear: strict minimisation at the left endpoint, strict minimisation at the right endpoint, strict minimisation at the centred position, and a symmetric double minimum at the two endpoints.
+In particular, the symmetric interior case with $(\beta_{0},\beta_{1})=(3.67,3.67)$ remains centred, while the symmetric endpoint case with $(\beta_{0},\beta_{1})=(0.2,0.2)$ keeps equal minima at both boundaries.
+$(\beta_{0},\beta_{1})=(0.2,2.8)$: left endpoint.
+$(\beta_{0},\beta_{1})=(2.8,0.2)$: right endpoint.
+$(\beta_{0},\beta_{1})=(3.67,3.67)$: unique interior point.
+$(\beta_{0},\beta_{1})=(0.2,0.2)$: two endpoint minimisers.
+Figure 2: Profiles of $a\mapsto\lambda_{q_{0}}(a)$ for the four canonical tests in Table 2. In each panel, the three curves correspond to $q_{0}=0.2$, $0.05$, and 0.
+For the main text, Table 2 together with Figure 2 gives a compact verification of Corollary 3.6.
+Additional diagnostic plots, such as the graph of $q_{0}\mapsto a^{\sharp}(q_{0})$, are useful during computation, but for the four canonical choices above they are less informative: the selector $a^{\sharp}(q_{0})$ is already constant on the sampled range of $q_{0}$.
+5 Conclusion
+We have studied the minimisation of the positive principal eigenvalue for an indefinite-weight problem under inhomogeneous Robin boundary conditions.
+After recalling the variational framework and the standard bang–bang reduction, we obtained a complete one-dimensional classification of the optimal favourable interval.
+The result shows that endpoint asymmetry in the Robin parameters leads to a genuinely richer picture than in the homogeneous case: depending on the balance between the two boundary parameters and the corresponding eigenvalue, the optimal interval may be left-attached, right-attached, or interior.
+This classification was then confirmed numerically by a transfer–matrix shooting method.
+We also introduced a Schrödinger-type extension in which the optimisable indefinite weight is coupled with a fixed nonnegative background potential.
+In the coercive setting, we proved that the principal-eigenvalue theory and the bang–bang structure persist.
+For the one-dimensional constant-potential family, we established a perturbative result showing that minimisers for small positive potential converge to those of the unperturbed problem, and we illustrated this behaviour numerically for representative configurations.
+These results are relevant from two complementary viewpoints.
+On the modelling side, they provide a mathematically explicit description of how favourable habitat should be placed when the two ends of the environment interact differently with the exterior and when an additional hostile background field is present.
+On the analytical side, they furnish a concrete bridge between indefinite-weight optimisation problems arising in reaction–diffusion equations and Schrödinger-type spectral optimisation.
+The transfer–matrix formulation also yields a reproducible computational framework that can be adapted to related one-dimensional optimisation problems.
+Several directions remain open.
+The most immediate one is to move beyond the perturbative constant-potential regime and seek a sharper description of the optimal placement for positive background potential.
+It would also be natural to investigate whether the present asymmetric Robin analysis can be combined with the change-of-variables techniques available for shells and other special geometries.
+More broadly, the higher-dimensional shape problem remains largely open, and the present one-dimensional results provide a useful benchmark for future analytical and numerical work.
+Acknowledgement
+This work was co-funded by the Czech Science Foundation (GAČR), Grant No. 25-16847S, and by the University of Ostrava, Grant No. SGS07/PŘF/2025.